# Channel Resolvability Using Multiplicative Weight Update Algorithm

### v3

@@ II Problem Formulation of channel resolvability
The problem is how to construct a codebook ${\cal C}$ such that for arbitrary small $\eta$, the total variation distance $d_{\mathrm{var}}(W_{P},W_{{\cal C}})$ is smaller than $\eta,$ as small size $L$ as possible.
@@ III Multiplicative weight update algorithm
MWU algorithm is an iterative algorithm that obtains approximately the best strategies. MWU algorithm has a few variations, and we use the so-called Hedge algorithm [5]. The MWU algorithm reviewed in Algorithm 1 is a classical version of the one presented in [22], and its performance is evaluated in the following lemma. For readers’ convenience, we provide a proof of the lemma in Appendix -A; see also [5, Theorem 2.3].
@@
  Fix $\varepsilon\in(0,\frac{1}{2}],$ and initialize the weight $w(1,y)=1$ for all $y\in{\cal Y}$. For each step $l=1,2,...,L$: 1. Calculate ${\cal Y}$-player’s mixed strategy as follows:
@@
2. ${\cal X}$-player selects a pure strategy $x_{l}\in{\cal X}$3. For the cost $(m(x_{l},y):y\in{\cal Y})$, update the weight as follows:
@@
(MWU algorithm performance) Fix an arbitrary $\varepsilon\in(0,\frac{1}{2}]$. The mixed strategies $f(1,\cdot),\ldots,f(L,\cdot)$ of ${\cal Y}$-player obtained by Algorithm 1 satisfies for every $y\in{\cal Y}$,
$(1-\varepsilon)\sum_{l=1}^{L}\bigg{(}\sum_{y^{\prime}\in{\cal Y}}f(l,y^{\prime})m(x_{l},y^{\prime})\bigg{)}\leq\sum_{l=1}^{L}m(x_{l},y)+\frac{\ln{|{\cal Y}|}}{\varepsilon}$
@@ IV Single-shot bound
In this section, we derive a single-shot bound for channel resolvability using the MWU algorithm. For the later convenience to apply the single-shot bound for the asymptotic setting, we consider truncated measures known as smoothing; e.g. see [36].
@@
(Truncated measures) For an arbitrary set ${\cal S}\subseteq{\cal X\times Y}$, and a given channel $W$, we define a truncated measure as
@@
and a truncated measure with input distribution $P$ as
@@
$D_{\mathrm{max}}$ ensures that the cost $m(x,y)$ is smaller than or equals to $1$ for all strategy $(x,y)\in{\cal X}\times\mathsf{s}(W_{P}^{\cal S})$. We run Algorithm 1 for $\mathsf{s}(W_{P}^{\cal S})$ instead of ${\cal Y}$; then, ${\cal Y}$-player’s mixed strategy $f(l,\cdot)$ is a distribution on $\mathsf{s}(W_{P}^{\cal S})$. Finally, we define ${\cal X}$-player’s strategy at each round $l$.
@@
For our choice of cost and ${\cal X}$-player’s strategy, we get the following performance of channel resolvability code ${\cal C}=\{x_{1},x_{2},\ldots,x_{L}\}$ obtained from ${\cal X}$-player’s strategies.
@@
Note that
$\sum_{y\in{\cal Y}}W_{P}^{{\cal S}^{c}}(y)=\sum_{(x,y)\in{\cal S}^{c}}P(x)W_{x}(y).$
Thus, we get Theorem 7 as long as the size $L$ satisfies assumption (3). ∎
@@ V Asymptotic result for fixed type
where $N(a,b|x^{n},y^{n})$ outputs the number of occurrences of set $(a,b)$ in $(x^{n},y^{n})$. We also define set
@@
where $\alpha^{\prime}=c_{W}\alpha$ for constant $c_{W}$ [15, Problem 2.5]. We also define the set ${\cal B}$ on the strong typical set of the output space ${\cal Y}^{n}$.
@@
:=${\cal B}:=\bigg{\{}y^{n}\in{\cal T}_{W_{T},\beta}^{n}:W_{P_{n}^{T}}(y^{n})\geq\frac{\tau}{|{\cal T}_{W_{T},\beta}^{n}|}\bigg{\}},$
where $\tau$ is a small positive value.
@@
To apply Theorem 7 for the set ${\cal S}$, we have to calculate $D_{\mathrm{max}}$, and $\sum_{(x^{n},y^{n})\in{\cal S}^{c}}P_{n}^{T}(x^{n})W_{x^{n}}(y^{n}).$ From Definition 3, we get
@@
$\leq-n(H(W|T)-\alpha^{\prime})+n(H(W_{T})+\beta^{\prime})-\ln{\tau}$
@@
where the inequality follows from equation (8) i.e.,
$-\frac{1}{n}\ln{W_{x^{n}}^{\cal S}(y^{n})}\geq H(W|T)-\alpha^{\prime},$
and the property of the set ${\cal B}$ with equation (9) i.e.,
$\begin{split}-\frac{1}{n}\ln{W_{P_{n}^{T}}^{\cal S}(y^{n})}&\leq-\frac{1}{n}\ln{\frac{\tau}{|{\cal T}_{W_{T},\beta}|}}\\ &\leq H(W_{T})+\beta^{\prime}-\frac{1}{n}\ln{\tau}.\end{split}$
Next, for ${\cal S}^{c}:=({\cal X}_{T}^{n}\times{\cal Y}^{n})\backslash{\cal S}$, we get
@@
Note also that the set ${\cal A}\cap\big{(}{\cal X}_{T}^{n}\times({\cal Y}^{n}\backslash{{\cal T}_{W_{T},\beta}^{n}})\big{)}$ is empty set, because for $\beta=|{\cal X}|\alpha$, each sequence $y^{n}$ in ${\cal A}$ is included in ${\cal T}_{W_{T},\beta}^{n}$ [15, Lemma 2.10]. Thus, we get
@@
$\qquad\qquad=\sum_{y^{n}\in{{\cal T}_{W_{T},\beta}^{n}}:\atop W_{P_{n}^{T}}(y^{n})<\frac{\tau}{|{\cal T}_{W_{T},\beta}^{n}|}}\sum_{x^{n}\in{\cal X}_{T}^{n}}P_{n}^{T}(x^{n})W_{x^{n}}(y^{n})$
$\qquad\qquad=\sum_{y^{n}\in{{\cal T}_{W_{T},\beta}^{n}}:\atop W_{P_{n}^{T}}(y^{n})<\frac{\tau}{|{\cal T}_{W_{T},\beta}^{n}|}}W_{P_{n}^{T}}(y^{n})$
$\qquad\qquad\leq\tau.$ (13)
Substituting equation (12) and (V) to equation (V), we get $\sum_{(x^{n},y^{n})\in{\cal S}^{c}}P_{n}^{T}(x^{n})W_{x^{n}}(y^{n})\leq\tau+\delta.$ Applying Theorem 7 for equation (V) and this equation, we get Theorem 10. ∎
@@ VI Asymptotic result for general input distribution
for each symbol $x^{n}\in{\cal X}^{n}$ where $P_{n}^{T}(x^{n}):=\frac{P_{n}(x^{n})}{P_{n}({\cal X}_{T}^{n})}\mathbf{1}[x^{n}\in{\cal X}_{T}^{n}]$. We first simulate the type distribution $P_{\cal T}$. For readers’ convenience, we provide a proof of the proposition in Appendix -B; This simulation is known as source resolvability, which is already proved by deterministic code construction [18, 35, 26]. However, we prove source resolvability using the MWU algorithm so that our channel resolvability code is constructed solely by the MWU algorithm.
@@
(Source resolvability with type) For type distribution $P_{\cal T}$, arbitrary small $\varepsilon^{\prime}\in(0,\frac{1}{2}],$ small positive $\tau^{\prime}>0$, and block length $n$, the codebook ${\cal C^{\prime}}=\{T_{1},T_{2},...,T_{L_{n}^{\prime}}\}\subset{\cal T}$ obtained by the MWU algorithm makes a distribution $R_{{\cal C}^{\prime}}(T)=\frac{1}{L_{n}^{\prime}}\sum_{l=1}^{L_{n}^{\prime}}\mathbf{1}[T_{l}=T]$. The distribution $R_{{\cal C}^{\prime}}(T)$ satisfies that $d_{\mathrm{var}}(P_{\cal T},R_{{\cal C}^{\prime}})\leq 2\varepsilon^{\prime}+\tau^{\prime},$ as long as the size $L_{n}^{\prime}$ satisfies
@@
(Channel resolvability for general input distribution) For a general input distribution $P_{n}$, a given i.i.d. channel $W,$ and small constants $\eta,\kappa$, the codebook ${\cal C}_{n}$ obtained by the MWU algorithm satisfies that $d_{\mathrm{var}}(W_{P_{n}},W_{{\cal C}_{n}})\leq\eta$ for sufficiently large $n$, as long as the size $L_{n}^{\prime\prime}$ satisfies that $L_{n}^{\prime\prime}\geq\exp(n(C(W)+\kappa)).$

### v4

@@ II Problem Formulation of channel resolvability
The problem "channel resolvability" is how to construct a codebook ${\cal C}$ such that for arbitrary small $\eta$, the total variation distance $d_{\mathrm{var}}(W_{P},W_{{\cal C}})$ is smaller than $\eta,$ as small size $L$ as possible.
@@ III Multiplicative weight update algorithm
The MWU algorithm is an iterative algorithm that obtains approximately the best strategies. The MWU algorithm has a few variations, and we use the so-called Hedge algorithm [5]. The MWU algorithm reviewed in Algorithm 1 is a classical version of the one presented in [22], and its performance is evaluated in the following lemma. For readers’ convenience, we provide a proof of the lemma in Appendix -A; see also [5, Theorem 2.3].
@@
 Fix $\varepsilon\in(0,\frac{1}{2}],$ and initialize weights $w(1,y)=1$ for $y\in{\cal Y}$. For each step $l=1,2,...,L$: 1. Calculate ${\cal Y}$-player’s mixed strategy as follows:
@@
2. ${\cal X}$-player selects a pure strategy $x_{l}\in{\cal X}$3. For the cost $(m(x_{l},y):y\in{\cal Y})$, weights are updated via:
@@
(MWU algorithm performance) Fix an arbitrary $\varepsilon\in(0,\frac{1}{2}]$. The mixed strategies $f(1,\cdot),\ldots,f(L,\cdot)$ of ${\cal Y}$-player’s strategy in Algorithm 1 satisfies for every $y\in{\cal Y}$,
$(1-\varepsilon)\sum_{l=1}^{L}\bigg{(}\sum_{y^{\prime}\in{\cal Y}}f(l,y^{\prime})m(x_{l},y^{\prime})\bigg{)}\leq\sum_{l=1}^{L}m(x_{l},y)+\frac{\ln{|{\cal Y}|}}{\varepsilon}.$
@@ IV Single-shot bound
In this section, we derive a single-shot bound for channel resolvability using the MWU algorithm. For the later convenience, we consider truncated measures known as smoothing, to apply the single-shot bound for the asymptotic setting; e.g. see [36].
@@
(Truncated measures) For an arbitrary set ${\cal S}\subseteq{\cal X\times Y}$, and a given channel $W$, we define truncated measure as
@@
and truncated measure with input distribution $P$ as
@@
$D_{\mathrm{max}}$ ensures that the cost $m(x,y)$ is smaller than or equals to 111 for all strategy $(x,y)\in{\cal X}\times\mathsf{s}(W_{P}^{\cal S})$. We run Algorithm 1 for $\mathsf{s}(W_{P}^{\cal S})$ instead of ${\cal Y}$; then, ${\cal Y}$-player’s mixed strategy $f(l,\cdot)$ is a distribution on $\mathsf{s}(W_{P}^{\cal S})$. Finally, we define ${\cal X}$-player’s strategy at each round $l$.
@@
For our choice of cost and ${\cal X}$-player’s strategy, we get the following performance of channel resolvability code ${\cal C}=\{x_{1},x_{2},\ldots,x_{L}\}$.
@@
Note that $\sum_{y\in{\cal Y}}W_{P}^{{\cal S}^{c}}(y)=\sum_{(x,y)\in{\cal S}^{c}}P(x)W_{x}(y).$ Thus, we get Theorem 7 as long as the size $L$ satisfies assumption (3). ∎
@@ V Asymptotic result for fixed type
[…]
@@
where $\alpha^{\prime}=c_{W}\alpha$ for constant $c_{W}$ [15, Problem 2.5]. We define the set ${\cal B}$ on the strong typical set of the output space ${\cal Y}^{n}$.
@@
:=${\cal B}:=\bigg{\{}y^{n}\in{\cal T}_{W_{T},\beta}^{n}:W_{P_{n}^{T}}^{\cal A}(y^{n})\geq\frac{\tau}{|{\cal T}_{W_{T},\beta}^{n}|}\bigg{\}},$
where $\tau$ is a small positive value and
$W_{P_{n}^{T}}^{\cal A}(y^{n})=\sum_{x^{n}\in{\cal X}^{n}}P_{n}^{T}(x^{n})W_{x^{n}}(y^{n})\mathbf{1}[(x^{n},y^{n})\in\mathcal{A}].$
@@
To apply Theorem 7 for the set ${\cal S}$, we have to calculate $D_{\mathrm{max}}$, and $\sum_{(x^{n},y^{n})\in{\cal S}^{c}}P_{n}^{T}(x^{n})W_{x^{n}}(y^{n}).$ First, we focus on the bound of $W_{x_{n}}^{\mathcal{S}}(y^{n})$ and $W_{P_{n}^{T}}^{\mathcal{S}}(y^{n})$. For $(x^{n},y^{n})\in\mathcal{S}$, equation (8) leads that
$\frac{1}{n}\ln{W_{x^{n}}^{\cal S}(y^{n})}\leq-H(W|T)+\alpha^{\prime},$ (10)
Note that
$W_{P_{n}^{T}}^{\mathcal{S}}(y^{n})=W_{P_{n}^{T}}^{\mathcal{A}}(y^{n}),$
for $y^{n}\in\mathsf{s}(W_{P_{n}^{T}}^{\mathcal{S}})\subset\mathcal{B}$. Using the property of the set ${\cal B}$ in equation (9), for $y^{n}\in\mathsf{s}(W_{P_{n}^{T}}^{\mathcal{S}})$, we obtain
$\frac{1}{n}\ln{W_{P_{n}^{T}}^{\cal S}(y^{n})}$ $=\ln{W_{P_{n}^{T}}^{\cal A}(y^{n})}$
$\geq\frac{1}{n}\ln{\frac{\tau}{|{\cal T}_{W_{T},\beta}^{n}|}}$
$\geq-H(W_{T})-\beta^{\prime}+\frac{1}{n}\ln{\tau}.$ (11)
Thus, from Definition 3, we get
@@
$\leq n(-H(W|T)+\alpha^{\prime})+n(H(W_{T})+\beta^{\prime})-\ln{\tau}$
@@
where the inequality follows from equation (10) and (V).
[…]
@@
Note also that the set ${\cal A}\cap\big{(}{\cal X}_{T}^{n}\times({\cal Y}^{n}\backslash{{\cal T}_{W_{T},\beta}^{n}})\big{)}$ is empty set, because each sequence $y^{n}$ in ${\cal A}$ is included in ${\cal T}_{W_{T},\beta}^{n}$ for $\beta=|{\cal X}|\alpha$ [15, Lemma 2.10]. We get
@@
$=\sum_{y^{n}\in{{\cal T}_{W_{T},\beta}^{n}}:\atop W_{P_{n}^{T}}^{\mathcal{A}}(y^{n})<\frac{\tau}{|{\cal T}_{W_{T},\beta}^{n}|}}\sum_{x^{n}\in{\cal X}_{T}^{n}}P_{n}^{T}(x^{n})W_{x^{n}}(y^{n})\mathbf{1}[(x^{n},y^{n})\in\mathcal{A}]$
$=\sum_{y^{n}\in{{\cal T}_{W_{T},\beta}^{n}}:\atop W_{P_{n}^{T}}^{\mathcal{A}}(y^{n})<\frac{\tau}{|{\cal T}_{W_{T},\beta}^{n}|}}W_{P_{n}^{T}}^{\mathcal{A}}(y^{n})\quad\leq\tau.$ (15)
Substituting equation (14) and (V) to equation (V), we get
$\sum_{(x^{n},y^{n})\in{\cal S}^{c}}P_{n}^{T}(x^{n})W_{x^{n}}(y^{n})\leq\tau+\delta.$
Applying Theorem 7 for equation (V) and this equation, we get Theorem 10. ∎
@@ VI Asymptotic result for general input distribution
for each symbol $x^{n}\in{\cal X}^{n}$ where
$P_{n}^{T}(x^{n}):=\frac{P_{n}(x^{n})}{P_{n}({\cal X}_{T}^{n})}\mathbf{1}[x^{n}\in{\cal X}_{T}^{n}].$
We first simulate the type distribution $P_{\cal T}$. For readers’ convenience, we provide a proof of the proposition in Appendix -B; This simulation is known as source resolvability, which is already proved by deterministic code construction [18, 35, 26]. However, we prove source resolvability using the MWU algorithm so that our channel resolvability code is constructed solely by the MWU algorithm.
@@
(Source resolvability with type) For type distribution $P_{\cal T}$, arbitrary small $\varepsilon^{\prime}\in(0,\frac{1}{2}],$ small positive $\tau^{\prime}>0$, and block length $n$, the codebook ${\cal C^{\prime}}=\{T_{1},T_{2},...,T_{L_{n}^{\prime}}\}\subset{\cal T}$ obtained by the MWU algorithm makes a distribution
$R_{{\cal C}^{\prime}}(T)=\frac{1}{L_{n}^{\prime}}\sum_{l=1}^{L_{n}^{\prime}}\mathbf{1}[T_{l}=T],$
such that
$d_{\mathrm{var}}(P_{\cal T},R_{{\cal C}^{\prime}})\leq 2\varepsilon^{\prime}+\tau^{\prime},$
as long as the size $L_{n}^{\prime}$ satisfies
@@
(Channel resolvability for general input distribution) For a general input distribution $P_{n}$, a given i.i.d. channel $W,$ and small constants $\eta,\kappa$, the codebook ${\cal C}_{n}$ obtained by the MWU algorithm satisfies that
$d_{\mathrm{var}}(W_{P_{n}},W_{{\cal C}_{n}})\leq\eta,$
for sufficiently large $n$, as long as the size $L_{n}^{\prime\prime}$ satisfies that
$L_{n}^{\prime\prime}\geq\exp(n(C(W)+\kappa)).$